OpenStudy (anonymous):

Are the number of permutations of two items from a data set always twice the number of combinations when taking two items from the same set?

OpenStudy (kropot72):

Yes. The number of permutations is two. The number of combinations is one.

OpenStudy (anonymous):

Yes. nPr = \[n! \over (n - 2)!\] nCr = \[n! \over (n - 2)! 2!\] When you simplify things, it will always be divided by 2 since 2! = 2 * 1 = 2.

OpenStudy (anonymous):

Awsome. I've been struggling with this one.